Jill wishes to provide herself, or her estate, with an income of $10,000 at the end of each year for 10 years. She will make a lump sum deposit when the account is established and add $3000 at the end of each year for 12 years. The income is to start at the end of the year following the year in which the last deposit was made. Compute the lump sum deposit. (All interest rates are 7 Percent Compounded annually)


1
Expert's answer
2021-09-05T17:12:02-0400

"\\dfrac{Po=d(1-(1 + \\frac{r}{k} \n\n\u200b\n ) -k.n}{\\frac{r}{k} }"

p= amount at the start

r= rate per period

n=number of periods

"\\dfrac{Po=10000(1-(1 + \\frac{0.07}{12} \n\n\u200b\n ) -12X10}{\\frac{0.07}{12} }"